In the cost function below, C(x) is the cost of producing x itemsl Find the average cost per item when the number of items is produced.
C(x)=4.3x+9,300
What is the average cost per item when 200 items are produced?
What is the average cost per item when 2000 items are produced?
What is the average cost per item when 5000 items are produced?
I am so confused.
Thank you so much.

Found 2 solutions by stanbon, flyoung:
Answer by stanbon(75887)

In the cost function below, C(x) is the cost of producing x items. Find the average cost per item when the number of items is produced.
C(x)=4.3x+9,300
What is the average cost per item when 200 items are produced?
C(200) = 4.3(200) + 9300 = $10,160
----------------------------------------
What is the average cost per item when 2000 items are produced?
C(2000) = 4.3(2000) + 9300 = $17,900
-----------------------------------------
What is the average cost per item when 5000 items are produced?
c(5000) = 4.3(5000)+9300 = $30,800
=========================================
Cheers,
Stan H.

Answer by flyoung(1)

The correct answer is: The way to find the AVERAGE cost per item is to use the cost equation and divide it by the number of items. For Example:
the average cost = C(x)/x = (4.3x + 9,300)/x.
So, if x = 200,
the equation should be average cost = AC(x)= (4.3*(200)+9,300)/200 = 50.8
if x=2000,
AC(x) = (4.3*2000+9300)/2000 = 8.95
If x=5000,
AC(x) = (4.3*5000+9300)/5000 = 6.16
Your question asks for the average cost PER ITEM, which is reasonable to assume that the cost would go down with more being made.
Total Cost = Variable cost + Fixed cost or TC = VC + FC
In your equation the fixed cost would be 9300. (set quantity x to zero and solve), while your variable cost would be 4.3x (the more you make the more it costs for supplies, etc).
The previous tutor StanH actually gave you the answer for total cost, NOT average cost.